OpenStudy (anonymous):

Which equation describes the line in slope-intercept form? slope = - 9 ; y-intercept= - 5/4

OpenStudy (bahrom7893):

y = -9x - 5/4

OpenStudy (bahrom7893):

y=Mx+b M - slope b - y-int

Directrix (directrix):

y = slope * x + y-intercept y = -9x - 5/4

OpenStudy (anonymous):

y= mx +c y = -9x -5/4